To precipitate all the sulfate ions from 22.0 mL of 0.150 molarity aluminum sulfate as barium sulfate, how many mL of 0.250 molarity3. barium nitrate are required? (8 points)



____Ba(NO3)2(aq)+____Al2(SO4)3(aq)→______BaSO4(s)+____Al(NO3)3(aq)



1
Expert's answer
2022-03-29T01:22:03-0400

2Ba(NO3)2 + Al2(SO4)3 = 3Ba(SO4) + 2Al(NO3)3

n(Al2(SO4)3) = M*V = 0.150*0.022 = 0.0033 mol

n(Ba(NO3)2) = 0.0033*3 = 0.0099 mol

V(Ba(NO3)2) = n/M = 0.0099/.0250 = 0.0396l (39.6 ml)


Answer 39.6ml